Help Please

Suppose you roll a six-sided die and flip three coins. What is the chance that the die will come up as an even number and you'll get at least one heads?

Express your answer as a value between 0 and 1, rounded to two decimal places
Answer #1

P(even number on die) = 3/6 = 1/2

P(at least one head) = 1 - P(no heads) = 1 - 1/8 = 7/8

Hence, P(even number on die and at least one heads) = (1/2)*(7/8) = 7/16 = 0.4375

So, the answer is 0.44 (rounded to 2 decimal places)

Please comment if any doubt. Thank you.